Why Range of particle will be maximum when dR/d alpha = 0?Pls show me it's proof with explanation.

Grade:11

1 Answers

Khimraj
5 years ago
Range R = u2sin2\alpha/2g
for maximum range dR/d\alpha = 0 {THis is the step forcalulating maxima of a function. You will read about this in mathematic differential calculus.}
So 2u2cos2\alpha/2g = 0
so \alpha = 45